How many grams of mercuric chloride must react to give 4.41 g of Hg?
Co(s)+HgCl2(aq)→CoCl3(aq)+Hg(l)
1
Expert's answer
2016-04-21T11:25:05-0400
2Co(s)+3HgCl2(aq)→2CoCl3(aq)+3Hg(l)
n (Hg) = n (HgCl2)
n = m/M
M (Hg) = 200.59 g/mol
M (HgCl2) = 271.5 g/mol
n (Hg)=4.41/200.59 = 0.02 mol
m (HgCl2) = n (HgCl2) · M (HgCl2) = 0.02 · 271.5 = 5.97 g